Suppose a consumer has income of $120 per period, and faces prices, price of X=2 and Price of Z =3. Her goal is to maximize her utility, described by the function U = 10X0.5Z0.5.


i. What is the consumer’s budget constraint? (3)


ii. State the formula for finding marginal utilities for goods X and Z (2)


iii. Calculate the marginal utilities for goods X and Z (4)


iv. State the utility maximizing condition for this consumer (4).


v. Calculate the utility maximizing bundle (X*, Z*) (13)

1
Expert's answer
2022-03-07T12:11:49-0500

i. Budget constraint is written as;

Y=PxX+PzZ

Where: Y=income

Px=Price of commodity X

Py=Price of commodity Y

120=2X+3Z........ budget constraint.

ii. Marginal utility(MU) is the change in utility with respect to a change quantity.

MUx="\dfrac{dU}{dX}"

MUy="\dfrac{dU}{dY}"

iii. MUx=5X-0.5Z0.5

MUy=5X0.5Z-0.5

iv. Utility is maximized when the slope of the budget line equals the slope of the utility function.

Slope of the budget line ="\dfrac{-P_x}{P_z}"

Slope of utility function ="\dfrac{-MU_x}{MU_z}"

Therefore, utility is maximized at;

"\dfrac{-P_x}{P_z}=\dfrac{-MU_x}{MU_z}"

v. Slope of the budget line ="\dfrac{-P_x}{P_z}" = "\dfrac{-2}{3}"

Slope of utility function ="\dfrac{-MU_x}{MU_z}=-\dfrac{5X^{-0.5}Z^{0.5}}{5X^{0.5}Z^{-0.5}}\\"

"=-\dfrac{Z^{0.5}Z^{0.5}}{X^{0.5}X^{0.5}}=\dfrac{-Z}{X}\\"

We have;

"\dfrac{-2}{3}=\dfrac{-Z}{X}"

Cross-multiplying, we have;

-2X=-3Z

Z="\dfrac{2}{3}X"

Inserting Z into the budget constraint, we have;

120=2X+3("\dfrac{2}{3}"X )

120=4X

X=30

Z="\dfrac{2}{3}" (30)

Z=20

Therefore, the optimizing bundle is;

(X,Z)=(30.20).